OpenStudy (anonymous):

One root of 5x^(2) +6x +k =0 is double the other. Find the roots and find the value of k

OpenStudy (anonymous):

you have to use the roots and coefficient rules of polynomials i.e. alpha + beta = -(b/a)

OpenStudy (cwrw238):

A + B = -6/5 and A*B = k / 5 let A = 2B so 2B + B = -6/5 2B*B = k/5 you need to solve this system for B and k

OpenStudy (anonymous):

i still dont get it